The kitchen floor of my home is 12'6" in width and 10'4" in length. The tiles I have selected are 5" on each side. I need to express my answer in improper fraction for the square footage. What is the area of each tile in square feet and how many tiles are required to tile the floor?
Answer by checkley77(12844)

5IN*IN=25IN^2 FOR EACH TILE
THERE ARE 12*12=144 IN^2 IN A 1 FT^2.
25/144 IS THE AREA OF EACH TILE IN FT^2.
(12 1/2)(10 1/3)=129 1/6 FT^2 OR (6*129+1)/6=(774+1)/6=775/6 FT^2 IS THE KITCHEN FLOOR AREA.
(775/6)/(25/144)=775/6*144/25=111,600/150 (~744 TILES WILL BE NEEDED).